OpenStudy (anonymous):

Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

sammixboo (sammixboo):

I will help you in just a sec! I am finishing up with another user

OpenStudy (anonymous):

okay thank you !

sammixboo (sammixboo):

Okay I am back! First plug in all the values. \(2 \times 2 ~-~2z4~+~ y2 ~- ~x2 ~+ ~z4\) \(2 \times 2 ~-~2\times 2 \times4~+~ 3\times2 ~- ~-4\times2 ~+ ~2\times4\) Now following the order of operations, we must multiply starting from the left side of the equation, so first what is 2 x 2?

OpenStudy (anonymous):

4

sammixboo (sammixboo):

Good now we have \(4~-~2\times 2 \times4~+~ 3\times2 ~- ~-4\times2 ~+ ~2\times4\) Now what is 2 x 2 x 4?

OpenStudy (anonymous):

16

sammixboo (sammixboo):

Great! now we have \(4 ~-~16~+~ 3\times2 ~- ~-4\times2 ~+ ~2\times4\) Now tell me what is 3 x 2?

OpenStudy (anonymous):

6

sammixboo (sammixboo):

Awesome now we have \(4 ~-~16~+~ 6 ~- ~(-4)\times2 ~+ ~2\times4\) Now can you tell me what is -4 x 3?

sammixboo (sammixboo):

Sorry -4 x 2, not 3

OpenStudy (anonymous):

-8

sammixboo (sammixboo):

Yes! Now we have \(4 ~-~16~+~ 6 ~- (-8)~+ ~2\times4\) Last multiplication :) What is 2 x 4?

OpenStudy (anonymous):

8

sammixboo (sammixboo):

Yes! Now we start subtracting/dding from left to right! \(4 ~-~16~+~ 6 ~- (-8)~+ ~8\) So what is 4 - 16?

OpenStudy (anonymous):

-12

sammixboo (sammixboo):

Good! Now we have \(-12~+~ 6 ~- (-8)~+ ~8\) Now what is -12 + 6?

OpenStudy (anonymous):

-6

sammixboo (sammixboo):

Correct! Now we have \((-6) ~- (-8)~+ ~8\) Can you tell me what is -6 - (-8)?

OpenStudy (anonymous):

2

sammixboo (sammixboo):

Yes! Now we have \(2~+ ~8\) :) And that is your answer

OpenStudy (anonymous):

10

sammixboo (sammixboo):

Yay!